The loudness of a heavy snore is 69dB. How many times as loud as a normal
conversation of 60dB is a heavy snore?

Answer by htmentor(1343)

Sound intensity in dB is given by dB = 10log10(I/I0), where I0 is a reference sound level. Since we will be taking ratios of two sound levels, we can make I0
any reference level we want. For simplicity let I0 = 1.
Then I1 = 10^(dB1/10) = 10^6, and I2 = 10^(dB2/10) = 10^6.9
I2/I1 = 7.943